OpenStudy (anonymous):

Please don't ignore.. but how do I simplify (4m+8/3n-3) / (2m+6/7n-7)?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=simplify+%284m%2B8%2F3n-3%29+%2F+%282m%2B6%2F7n-7%29 That is as simple as it gets (see the link above)

OpenStudy (anonymous):

It looks like this: 4m + 8 2m + 6 ------ / ------- 3n - 3 7n-7

OpenStudy (jamesj):

First \[ (4m+8/3n-3) / (2m+6/7n-7) = \frac{4m+8}{3n-3} \frac{7n-7}{2m+6} \]

OpenStudy (anonymous):

I did that but I'm stuck..

OpenStudy (jamesj):

Now you factor out lots of constants: for instance 4m + 8 = 4(m + 2) and 2m + 6 = 2(m + 3) 3n - 3 = 3(n - 1) 7n = 7 = 7(n - 1) Once you do that, you see that the n-1 terms cancel and you can simplify.

OpenStudy (anonymous):

yup. I got 4(m+2) x 7 --------- 3x2(m-3

OpenStudy (jamesj):

Keep simplifying, cancel out a 2 from top and bottom (and note your sign error on the bottom: m + 3, not m - 3)

OpenStudy (anonymous):

ohhh! & then I got: 2(m+2)x7 -------- 3(m+3) Do I just get: 2m + 4 x 7 --------- 3m+3 ?

OpenStudy (jamesj):

Keep the terms factorized. I'd write it as: \[ \frac{14 } {3 } . \frac{m+2}{m+3} \]

OpenStudy (anonymous):

ohhhh! OMG, thank you so much!!

OpenStudy (jamesj):

Notice by the way in what did your last step you didn't handle the x7 very well. It now looks as though it 2 x 7 when it should be the whole expression x7

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

but the whole expression "x7"?

OpenStudy (jamesj):

You have a 7 in the numerator, but what you wrote: 2m + 4 x 7 --------- 3m+3 makes it look as though it's just 7 times the 4; not 7 x (2m + 4) or (2m + 4) x 7 which is what it should be. So I'm just be careful with the way you manipulate these expressions, as you've just introduced an error.